OpenStudy (anonymous):

Solve x + 7 = –2(4x + 10). You must show all work

OpenStudy (anonymous):

@hero can you help you helped me once this week and your the only one i understood

OpenStudy (anonymous):

@ksaimouli can you show me step by step please

OpenStudy (ksaimouli):

multiply -2 to 4x+10 what do u get

jimthompson5910 (jim_thompson5910):

Similar Problem (same idea, the numbers are changed) You want to isolate x. Here's how you do it: x+3 = -5(2x+6) ... Start with the given equation. x+3 = -10x-30 ... Distribute x = -10x-30-3 ... Subtract 3 from both sides. x+10x = -30-3 ... Add 10x to both sides. 11x = -33 ... Combine like terms. x = -33/11 ... Divide both sides by 11 to fully isolate x. x = -3 ... Divide So the final answer is x = -3 Note: this is NOT the answer to your problem since it's only a similar example.